Ezekiel manages a clothing store. In May, sales were $129,800. In June, sales were $131,975. What was the rate of change?
Answer #1

Rate of change = (Current sales / Previous sales - 1) * 100

Given: Current sales = $131975, previous sales = $129800

Putting these values in the above formula, we get,

Rate of change = ($131975 / $129800) - 1) * 100

Rate of change = (1.0167565 -1) * 100

Rate of change = 0.0167565 * 100 = 1.675%